# Transforming Time-Varying to Static Channels: The Power of Fluid Antenna Mobility

### v1

@@ II Channel Model
$\begin{array}[]{l}{\beta_{p}^{{\rm{NLoS}}}}=\sqrt{\frac{{{K_{s}}^{\prime}}}{{{P_{s}}\sum\limits_{s=1}^{S}{{K_{s}}^{\prime}}}}}.\end{array}\!$ (8)
@@ III-C The port selection method for a LoS channel
$\begin{array}[]{l}{m(\Delta t)}=\mathop{\arg\min}\limits_{{m(\Delta t)}}({\sin^{2}}{\varsigma^{{\rm{LoS}}}}).\end{array}\!$ (31)
@@ III-D The port selection method for a multi-path channel
$\begin{array}[]{l}({f}(x))_{\min}=\min\left(f(0),f({{W\lambda}})\right).\end{array}\!$ (48)
@@ V Numerical Results
Delay spread (ns) 616
CSI delay (ms) 4
UEs speed (km/h) 60, 120
@@ VI Conclusion
In this paper, we address the mobility problem with the assistance of FA. We propose a matrix pencil-based moving port prediction method, which transforms the time-varying channel to the static channel by sliding the liquid. We propose a port selection method to obtain the optimal port. In the theoretical analysis, we derive the minimum FA length and the maximum sliding speed. For a LoS channel, we also prove that the prediction MSE converges to zero if the density of ports and the number of BS antennas are large enough. For a multi-path channel, we also derive the closed-form lower and upper bounds of the prediction MSE. Simulation results demonstrate that, with the assistance of FA, our proposed MPMP method provides significant gain in the high-mobility scenario.
@@
$\begin{array}[]{l}\!\!\!\!{v_{f}}\!=\!\frac{{d_{v}^{\rm{rx}}}{\left|{{m}({t_{0}+T})-{m}({t_{0}})}\right|}}{{T}}\!\approx\!\left|{\left[{\frac{{\lambda\omega^{{\rm{LoS}}}}}{{\cos\theta_{\rm{EOA}}^{{\rm{LoS}}}}}+\frac{{\lambda({k({{t_{0}}+T})}-{k({{t_{0}}})})}}{T{\cos\theta_{\rm{EOA}}^{{\rm{LoS}}}}}}\right]}\right|.\end{array}\!\!\!\!\!\!$ (72)
@@
$\begin{array}[]{l}y\in(-0.5,0.5].\end{array}\!$ (80)
@@
$\begin{array}[]{l}\mathop{\lim}\limits_{\rho\to\infty}{\left\|{{\varepsilon_{n}}}\right\|_{2}^{2}}_{\max}=\mathop{\lim}\limits_{\rho\to\infty}\frac{{{K_{R}}}}{{{K_{R}}+1}}(2-2\cos(\frac{{\pi}}{\rho}))=0.\end{array}\!$ (81)
@@
with ${k_{p}}=2\pi f{\tau_{p}}-{\delta^{{\rm{LoS}}}}-{\varsigma^{{\rm{LoS}}}}$, and

### v2

@@ II Channel Model
$\begin{array}[]{l}{\beta_{p}^{{\rm{NLoS}}}}=\sqrt{\frac{{{K_{s}}^{\prime}}}{{{P_{s}}\sum\limits_{s=1}^{S}{{K_{s}}^{\prime}}}}}\end{array}.\!$ (8)
@@ III-C The port selection method for a LoS channel
$\begin{array}[]{l}{m(\Delta t)}=\mathop{\arg\min}\limits_{{m(\Delta t)}}({\sin^{2}}{\varsigma^{{\rm{LoS}}}})\end{array}.\!$ (31)
@@ III-D The port selection method for a multi-path channel
$\begin{array}[]{l}({f}(x))_{\min}=\min\left(f(0),f({{W\lambda}})\right)\end{array}.\!$ (48)
@@ V Numerical Results
Delay spread ($\rm{ns}$) 616
CSI delay ($\rm{ms}$) 4
UEs speed (${\rm{km/h}}$) 60, 120
@@ VI Conclusion
In this paper, we addressed the mobility problem with the assistance of FA. We proposed a matrix pencil-based moving port prediction method, which transforms the time-varying channel to the static channel by sliding the liquid. We proposed a port selection method to obtain the optimal port. In the theoretical analysis, we derived the minimum FA length and the maximum sliding speed. For a LoS channel, we also proved that the prediction MSE converges to zero if the density of ports and the number of BS antennas are large enough. For a multi-path channel, we also derived the closed-form lower and upper bounds of the prediction MSE. Simulation results demonstrate that, with the assistance of FA, our proposed MPMP method provides significant gain in the high-mobility scenario.
@@
$\begin{array}[]{l}{v_{f}}\!=\!\frac{{d_{v}^{\rm{rx}}}{\left|{{m}({t_{0}+T})-{m}({t_{0}})}\right|}}{{T}}\!\approx\!\left|{\left[{\frac{{\lambda\omega^{{\rm{LoS}}}}}{{\cos\theta_{\rm{EOA}}^{{\rm{LoS}}}}}+\frac{{\lambda({k({{t_{0}}+T})}-{k({{t_{0}}})})}}{T{\cos\theta_{\rm{EOA}}^{{\rm{LoS}}}}}}\right]}\right|\end{array}.$ (72)
@@
$\begin{array}[]{l}y\in(-0.5,0.5]\end{array}.\!$ (80)
@@
$\begin{array}[]{l}\mathop{\lim}\limits_{\rho\to\infty}{\left\|{{\varepsilon_{n}}}\right\|_{2}^{2}}_{\max}=\mathop{\lim}\limits_{\rho\to\infty}\frac{{{K_{R}}}}{{{K_{R}}+1}}(2-2\cos(\frac{{\pi}}{\rho}))=0\end{array}.\!$ (81)
@@
with ${k_{p}}=2\pi f{\tau_{p}}-{\delta^{{\rm{LoS}}}}-{\varsigma^{{\rm{LoS}}}}$, ${b_{p}}=2{D_{n,h}}\sin{\theta_{p}}\sin{\phi_{p}}+2{D_{n,v}}\cos{\theta_{p}}$, and